THE AI OF PHILOSOPHY John McCarthy Stanford University jmc@cs.stanford.edu http://www-formal.stanford.edu/jmc/

• A slogan for AI: Whatever a person can do, he should be able to make a computer do for him.

• AI is the continuation of philosophy by other means. – van Benthem

• AI, as you’ll see, takes sides in long standing controversies.

• What concepts and advice does AI have to offer philosophy?

CONCEPTS AND ADVICE

• Start small. Beliefs of a thermostat. Extend to partly known temperature control system.

• Dennett: intentional stance, design stance → McC: designer’s stance.

• Axiomatizations of the effects of actions and other events. Situation calculus.

• Formalized nonmonotonic reasoning. Predicate completion and circumscription.

MORE CONCEPTS AND ADVICE

• Elaboration tolerance – One of the missionaries is Jesus Christ. English is still better than present AI formalisms but relies on human common sense.

• Approximate objects and theories. Mount Everest, purposes of the United States.

• Concepts as objects. “I thought your yacht was longer than it is”. ConceptP erson(I, LLength(Y Y ourY acht)) > Length(Y ourY acht)

COMMON SENSE INFORMATIC SITUATION

• Bounded theories have to be revised from the outside. Human-level common sense theories must be extendable as we humans extend our own theories.

• open ended theories, the theory as an object, nonmonotonic reasoning, elaboration tolerance, approximate entities.

CONTEXTS AS OBJECTS

• Ist(Context-of-this-lecture(Speaker = McCarthy))

• Relations among contexts, lifting rules. Ist(Outer, V alue(AF -context, P rice(engine)) = V alue(GE-context(P rice(engine)))

+V alue(GE-context(P rice(Spare-parts-kit(engine))))). (1)